OpenStudy (anonymous):

given a point (-1, -3) on the terminal side of theta. Find the exact value of the six trigonometric functions of theta.

OpenStudy (turingtest):

 is this the correct figure?

OpenStudy (anonymous):

\[\text{Sin}[\theta ]=\frac{3}{\sqrt{10}},\text{Cos}[\theta ]=\frac{1}{\sqrt{10}},\text{Tan}[\theta ]=3,\text{Cot}[\theta ]=\frac{1}{3}, \]\[\text{Sec}[\theta ]=\sqrt{10},\text{Csc}[\theta ]=\frac{\sqrt{10}}{3} \]